Average Rate of Return Method, Net Present Value Method, and Analysis

The capital investment committee of Ellis Transport and Storage Inc. is considering two investment projects. The estimated income from operations and net cash flows from each investment are as follows:

Warehouse Tracking Technology
Year Income from
Operations
Net Cash
Flow
Income from
Operations
Net Cash
Flow
1 $52,200 $161,000 $110,000 $258,000
2 52,200 161,000 84,000 217,000
3 52,200 161,000 42,000 153,000
4 52,200 161,000 18,000 105,000
5 52,200 161,000 7,000 72,000
Total $261,000 $805,000 $261,000 $805,000

Each project requires an investment of $580,000. Straight-line depreciation will be used, and no residual value is expected. The committee has selected a rate of 10% for purposes of the net present value analysis.

Present Value of $1 at Compound Interest
Year 6% 10% 12% 15% 20%
1 0.943 0.909 0.893 0.870 0.833
2 0.890 0.826 0.797 0.756 0.694
3 0.840 0.751 0.712 0.658 0.579
4 0.792 0.683 0.636 0.572 0.482
5 0.747 0.621 0.567 0.497 0.402
6 0.705 0.564 0.507 0.432 0.335
7 0.665 0.513 0.452 0.376 0.279
8 0.627 0.467 0.404 0.327 0.233
9 0.592 0.424 0.361 0.284 0.194
10 0.558 0.386 0.322 0.247 0.162

Required:

1a. Compute the average rate of return for each investment. If required, round your answer to one decimal place.

Average Rate of Return
Warehouse %
Tracking Technology %

1b. Compute the net present value for each investment. Use the present value of $1 table above. If required, use the minus sign to indicate a negative net present value.

Warehouse Tracking Technology
Present value of net cash flow total $ $
Less amount to be invested $ $
Net present value $ $

2. The warehouse has a   net present value as tracking technology cash flows occur   in time. Thus, if only one of the two projects can be accepted, the   would be the more attractive.
Answer #1

Solution 1a:

Average rate of return = Average annual income / Average investment

Average annual income:

Warehouse = $261,000 / 5 = $52,200

Tracking technology = $261,000 / 5 = $52,200

Average investment:

Warehouse = ($580,000 + 0) / 2 = $290,000

Tracking technology =  ($580,000 + 0) / 2 = $290,000

Average rate of return:

Warehouse = $52,200 / $290,000 = 18%

Tracking technology = $52,200 / $290,000 = 18%

Solution 1b:

Computation of NPV
Warehouse Tracking Technology
Particulars Period PV Factor Amount Present Value Amount Present Value
Cash outflows:
Cost of Equipment 0 1 $580,000 $580,000 $580,000 $580,000
Present Value of Cash outflows (A) $580,000 $580,000
Cash Inflows
Year 1 1 0.909 $161,000 $146,349 $258,000 $234,522
Year 2 2 0.826 $161,000 $132,986 $217,000 $179,242
Year 3 3 0.751 $161,000 $120,911 $153,000 $114,903
Year 4 4 0.683 $161,000 $109,963 $105,000 $71,715
Year 5 5 0.621 $161,000 $99,981 $72,000 $44,712
Present Value of Cash Inflows (B) $610,190 $645,094
Net Present Value (NPV) (B-A) $30,190 $65,094

Solution 2:

The Warehouse has a lower net present value as tracking technology cash flow occur earlier in time. Thus, if only one of the two projects can be accepted, the tracking technology would be the more attractive.